# Transcendence and normality of complex numbers via Hurwitz continued fractions

### v2

@@ 7.1. Preliminary lemmas
Take $n,k\in\mathbb{N}$, $(x_{1},\ldots,x_{n})\in\mathscr{D}^{n}$, $(d_{1},\ldots,d_{k})\in\mathscr{D}^{k}$, $m\in\mathbb{Z}$ such that $|m|\geq 2$ and $c\in\{m,im\}$. If $(x_{1},\ldots,x_{n},c,d_{1},\ldots,d_{k})\in\mathsf{R}(n+k+1)$, then
@@
For $k=0$, we have $\mathfrak{F}^{\circ}_{n+1}(x_{1},\ldots,x_{n},c)=\mathfrak{F}^{\circ}(c)$.
@@
When $k=0$, the lemma can be shown by checking case by case all the options for $\mathfrak{F}^{\circ}_{n}(x_{1},\ldots,x_{n})$ and $c$. For example, if $\mathfrak{F}^{\circ}_{n}(x_{1},\ldots,x_{n})=\mathfrak{F}^{\circ}_{1}(2)$, then $c=im$ for some $m\in\mathbb{Z}$ with $|m|\geq 2$ or $c=m$ for some $m\in\mathbb{N}_{\geq 2}$ (see Figure 5); in either case, $\mathfrak{F}^{\circ}_{n+1}(x_{1},\ldots,x_{n},c)=\mathfrak{F}^{\circ}(c)$. Assume that $k\geq 1$. If $|c|\geq 3$, then $\mathfrak{F}^{\circ}_{n+1}(x_{1},\ldots,x_{n},c)=\mathfrak{F}^{\circ}$ and, by (8),
@@
$=\mathfrak{F}^{\circ}_{k}(d_{1},\ldots,d_{k}).$
When $|c|=2$, we check the $4$ possible forms of $\mathfrak{F}^{\circ}_{n+1}(x_{1},\ldots,x_{n},c)\neq\varnothing$ to conclude that if $(x_{1},\ldots,x_{n},c,d_{1})\in\mathsf{R}(n+2)$ then
$\mathfrak{F}^{\circ}_{n+2}(x_{1},\ldots,x_{n},c,d_{1})=\mathfrak{F}^{\circ}_{1}(d_{1}).$
@@
$\iota[\mathfrak{F}^{\circ}_{n+1}(x_{1},\ldots,x_{n},2)]=\iota\left[\mathfrak{F}^{\circ}_{1}(2)\right]=\iota[\mathfrak{F}^{\circ}]\cap\left\{z\in\mathbb{C}:\operatorname{Re}(z)>\frac{1}{2}\right\},$
so, for all $d\in\mathscr{D}$, we have
$\tau_{d}[\mathfrak{F}^{\circ}]\cap\iota[\mathfrak{F}^{\circ}_{n+1}(x_{1},\ldots,x_{n},c)]=\varnothing\;\text{ or }\;\tau_{d}[\mathfrak{F}^{\circ}]\cap\iota[\mathfrak{F}^{\circ}]$
(cfr. Figures 1 and 4). From here, we argue as in the case $|c|\geq 3$. ∎
Lemma 7.2 follows from the definitions of $\mathsf{mir}_{1}(z)=\overline{z}$, $\mathsf{mir}_{2}(z)=-\overline{z}$, and $S$. Lemma 7.3 is a corollary of Lemma 3.5 and its proof.
@@
If $m\in\mathbb{Z}$ is such that $|m|\geq 2$ and $a\in\{1+im,m+i\}$, then
@@
Take $\mathbf{b}=(b_{n})_{n\geq 1}\in\Omega$. If $j\in\mathbb{N}$ is such that $(b_{1},\ldots,b_{j})\in\mathsf{R}(j)$ but $(b_{1},\ldots,b_{j})\in\mathsf{Ir}(j+1)$, then
@@
Lemma 7.4.
If $\mathsf{mir}\in\mathsf{Mir}$, then $\overline{\Lambda}\circ\mathsf{mir}=\mathsf{mir}\circ\overline{\Lambda}$.
@@
The lemma follows from the continuity of $\mathsf{mir}_{1}$ and $\mathsf{mir}_{2}$ from the fact that for every $\mathsf{mir}\in\mathsf{Mir}$ we have
$\mathsf{mir}\left([0;z_{1},\ldots,z_{n}]\right)=[0;\mathsf{mir}(z_{1}),\ldots,\mathsf{mir}(z_{n})]$
for any $z_{1},\ldots,z_{n}\in\mathbb{C}$ such that both sides of the equality are defined. ∎
7.2. Proof of the algorithm
Let $\mathbf{a}=(a_{n})_{n\geq 1}$ be an irregular sequence. Since every cylinder of level $1$ is regular, there is some $j_{0}\in\mathbb{N}$ such that
@@ 7.2. Proof of the algorithm
Note that $\mathfrak{F}^{\circ}_{j_{0}}(a_{1},\ldots,a_{j_{0}})\neq\mathfrak{F}^{\circ}$, otherwise $(a_{1},\ldots,a_{j_{0}},a_{j_{0}+1})$ would be regular. The open prototype set $\mathfrak{F}^{\circ}_{j_{0}}(a_{1},\ldots,a_{j_{0}})$ can assume twelve forms (see Figure 1); by the proof of Lemma 3.5, we may conclude the existence of some $m\in\mathbb{Z}$ such that
@@
Assume that $|m|\geq 3$, then
$\mathfrak{F}^{\circ}_{j_{0}+1}\left(a_{1},\ldots,a_{j_{0}},S(a_{j_{0}+1})\right)=\mathfrak{F}^{\circ}\quad\text{ and }\quad(b_{1}(1),\ldots,b_{1}(j_{0}+1))\in\mathsf{R}(j_{0}+1).$
If $a_{j_{0}+1}=1+im$, by Lemma 7.3 and $[0;a_{j_{0}+2},a_{j_{0}+3},\ldots]\in\mathfrak{F}$, we have
@@
$=[1-\overline{a_{j_{0}+1}};\overline{a_{j_{0}+2}},\overline{a_{j_{0}+3}},\ldots]$
@@
$=[0;b_{1}(1),b_{1}(2),\ldots,b_{1}(j_{0}),b_{1}(j_{0}+1),b_{1}(j_{0}+2),b_{1}(j_{0}+3),\ldots]=\overline{\Lambda}(\mathbf{b}_{1}).$
Similarly, if $a_{j_{0}+1}=m+i$, then Lemma 7.3 and $[0;a_{j_{0}+2},a_{j_{0}+3},\ldots]\in\mathfrak{F}$ give
@@
Then, since $z=\overline{z}+i$ whenever $\operatorname{Im}(z)=\frac{1}{2}$, we have
@@
$=[S\left(a_{j_{0}+1}\right);\mathsf{mir}_{1}(a_{j_{0}+1}),\mathsf{mir}_{1}(a_{j_{0}+2}),\mathsf{mir}_{1}(a_{j_{0}+3}),\ldots].$
We conclude $\Lambda(\mathbf{a})=\overline{\Lambda}(\mathbf{b}_{1})$.
Assume that $|m|=2$. Then, applying $\iota$ on the sets in Lemma 3.5, we can see that $a_{j_{0}+1}\in\{1+2i,1-2i,2+i,-2+i\}$. Suppose that $a_{j_{0}+1}=1+2i$, then we must have either
@@
(to see it, apply $\iota$ on every set in Lemma 3.5 and notice that the only ones leading to an irregular cylinder after concatenating $1+2i$ are $\mathfrak{F}^{\circ}_{1}(-1-i)$ and $\mathfrak{F}^{\circ}_{1}(-2)$). In either case, we have
$\mathfrak{F}^{\circ}_{j_{0}+1}\left(a_{1},\ldots,a_{j_{0}},S(a_{j_{0}+1})\right)=\mathfrak{F}^{\circ}_{j_{0}+1}(a_{1},\ldots,a_{j_{0}},2i)=\mathfrak{F}^{\circ}_{1}(2i)\neq\varnothing.$
A similar argument holds when $a_{j_{0}+1}\in\{1-2i,2+i,2-i\}$. Irrespective to what case occurs when $|m|=2$, we may show $\Lambda(\mathbf{a})=\overline{\Lambda}(\mathbf{b}_{1})$ as when $|m|\geq 3$. To sum up, the sequence $\mathbf{b}_{1}$ has the next properties:
  1. i.
    $(b_{1}(1),\ldots,b_{1}(j_{0}+1))\in\mathsf{R}(j_{0}+1)$,
  2. ii.
    $\Lambda(\mathbf{a})=\overline{\Lambda}(\mathbf{b}_{1})$.
Assume that for some $N\in\mathbb{N}$ we have we are able to perform $N$ iterations of the algorithm and the resulting sequence $\mathbf{b}_{N}$ satisfies:
  1. i.
    $(b_{N}(1),\ldots,b_{N}(j_{N-1}+1))\in\mathsf{R}(j_{N-1}+1)$,
  2. ii.
    $\Lambda(\mathbf{a})=\overline{\Lambda}(\mathbf{b}_{N})$.
  3. iii.
    Either
    (25) $\operatorname{Re}\left([a_{j_{N-1}+1};a_{j_{N-1}+2},\ldots]_{\mathbb{C}}\right)=\frac{1}{2}\;\text{ or }\;\operatorname{Im}\left([a_{j_{N-1}+1};a_{j_{N-1}+2},\ldots]_{\mathbb{C}}\right)=\frac{1}{2}.$
@@
Next, we discuss the $(N+1)$-th iteration. Suppose there is some $j_{N}\in\mathbb{N}$ satisfying
(26) $\left(b_{N}(1),\ldots,b_{N}(j_{N})\right)\in\mathsf{R}(j_{N})\quad\text{ and }\quad\left(b_{N}(1),\ldots,b_{N}(j_{N}+1)\right)\not\in\mathsf{R}(j_{N}+1).$
Let us show that
@@
(29) $\Lambda(\mathbf{a})=\overline{\Lambda}(\mathbf{b}_{N+1}).$
Note that $j_{N}-j_{N-1}\in\mathbb{N}$ by the induction hypothesis. We consider separately the cases $j_{N}-j_{N-1}=1$ and $j_{N}-j_{N-1}\geq 2$.
Case $j_{N}=j_{N-1}+1$. Since $b_{N}(j_{N})=S(b_{N-1}(j_{N}))=S(b_{N-1}(j_{N-1}+1))$, there is some $m\in\mathbb{Z}$, $|m|\geq 2$, such that $b_{N}(j_{N})\in\{m,im,-m,-im\}$. If we had $|m|\geq 3$, by Lemma 7.1 we would have
@@
which implies $(b_{N}(1),\ldots,b_{N}(j_{N}),b_{N}(j_{N}+1))\in\mathsf{R}(j_{N}+1)$, contradicting (26). Therefore, $|m|=2$ and $b_{N}(j_{N})\in\{-2,-2i,2,2i\}$. Let us further assume that $b_{N}(j_{N})=-2$, the other possibilities are treated similarly. Then, we have
@@
Suppose that $b_{N-1}(j_{N})=-2+i$, an analogous argument works for $-2-i$. Under these new assumptions, Condition (26) and Lemma 7.1 imply
(30) $\mathfrak{F}^{\circ}_{j_{N}}\left(b_{N}(1),\ldots,b_{N}(j_{N})\right)=\mathfrak{F}^{\circ}_{1}(-2)\quad\text{ and }\quad\operatorname{Re}\left(b_{N}(j_{N}+1)\right)\geq 1$
(observe that $\mathfrak{F}^{\circ}_{1}(-2)=-\mathfrak{F}^{\circ}_{1}(2)$, $\iota[\mathfrak{F}^{\circ}_{1}(-2)]=-\iota[\mathfrak{F}(2)]$ and see Figure 5). We claim that
@@
Certainly, when $N=1$, we may consider $\mathsf{mir}$ to be the identity map on $\mathbb{C}$. For $N>1$, the induction hypothesis gives $j_{N-1}\geq j_{N-2}+1$, so $j_{N-1}+n\geq j_{N-2}+n+1$ for all $n\in\mathbb{N}$ and (31) follows from Remark 7.5.
Next, we consider every possibility for $\mathsf{mir}$ in (31) separately. Recall that
@@
  • When $\mathsf{mir}$ is the identity map, we have $a_{j_{N}}=-2+i$ and $\operatorname{Re}(a_{j_{N}+1})\geq 1$. Hence, the set $\mathfrak{F}_{j_{N}}(a_{1},\ldots,a_{j_{N}})\in\mathsf{Ir}(j_{N})$ is of one of the forms depicted on Figure 8 (this can be checked inductively using (3)). We consider each case independently.
    1. (a)
      If $\mathfrak{F}_{j_{N}}(a_{1},\ldots,a_{j_{N}})=\{[0;a_{j_{N}+1},a_{j_{N}+2},\ldots]_{\mathbb{C}}\}=\{\zeta_{4}\}$, we conclude (28) from $\operatorname{Re}(\zeta_{4}^{-1})=\frac{1}{2}$. Moreover, we have
@@
      $\overline{\Lambda}\left(b_{N}(j_{N}),b_{N}(j_{N}+1),\ldots\right)=\overline{\Lambda}\left(b_{N+1}(j_{N}),b_{N+1}(j_{N}+1),\ldots\right)$
      and, since $b_{N}(k)=b_{N+1}(k)$ for all $k\in\{1,\ldots,j_{N}-1\}$, Equation (29) follows.
    2. (b)
      Suppose that $\mathfrak{F}_{j_{N}}(a_{1},\ldots,a_{j_{N}})$ is the arch of $C(1-i)$ going from $-\zeta_{1}=\frac{1}{2}-i\alpha$ (excluded) to $\zeta_{4}$ (included). Then,
@@
      so $\zeta_{2}=[0;a_{j_{N+1}},a_{j_{N+2}},\ldots]$ and we argue as in Case (a).
    3. (c)
      When $\mathfrak{F}_{j_{N}}(a_{1},\ldots,a_{j_{N}})$ is the arch of $C(1)$ from $\overline{\zeta_{4}}$ (excluded) to $\zeta_{4}$ (included), we argue as in case (b).
    4. (d)
@@
      and $\zeta_{2}=[0;a_{j_{N}+1},a_{j_{N}+2},\ldots]_{\mathbb{C}}$. Now, we proceed as in case (a).
@@
  • If $\mathsf{mir}=\mathsf{mir}_{1}$, then $a_{j_{N}}=-2-i$ but (25) cannot hold.
  • When $\mathsf{mir}=\mathsf{mir}_{2}$, then $a_{j_{N}}=2+i$, $\operatorname{Re}(a_{j_{N}+1})\leq-1$ and, by (25),
    (32) $\operatorname{Im}([a_{j_{N}};a_{j_{N}+1},a_{j_{N}+2},\ldots]=\frac{1}{2}.$
@@
    1. (a)
      If $\mathfrak{F}_{j_{N}}(a_{1},\ldots,a_{j_{N}})=\{\zeta_{2}\}=\{[0;a_{j_{N}+1},a_{j_{N}+2},\ldots]\}$, then $[a_{j_{N}};a_{j_{N}+1},\ldots]=-2+i+\zeta_{2}$, contradicting (32).
    2. (b)
      If $\mathfrak{F}_{j_{N}}(a_{1},\ldots,a_{j_{N}})$ were the arch of $C(-1-i)$ going from $\zeta_{3}$ (included) to $\zeta_{2}$ (included), by (25) we would obtain $\zeta_{3}=[0;a_{j_{N}+1},a_{j_{N}+2},\ldots]_{\mathbb{C}}$, but $a_{j_{N}+1}=2i$ contradicts $\operatorname{Re}(a_{j_{N}+1})\leq-1$.
    3. (c)
      The impossibility of $\mathfrak{F}_{j_{N}}(a_{1},\ldots,a_{j_{N}})$ being the arch of $C(-1)$ from $\zeta_{2}$ (included) to $\overline{\zeta_{2}}$ (excluded) is shown as in case (b).
    4. (d)

### v3

@@ 7.1. Auxiliary lemmas
Take $n\in\mathbb{N}$, $(x_{1},\ldots,x_{n})\in\mathscr{D}^{n}$, $m\in\mathbb{Z}$ with $|m|\geq 2$, and $c\in\{m,im\}$.
  1. i.
    If $(x_{1},\ldots,x_{n},c)\in\mathsf{R}(n+1)$, then $\mathfrak{F}^{\circ}_{n+1}(x_{1},\ldots,x_{n},c)=\mathfrak{F}^{\circ}_{n}(c)$.
  2. ii.
    For all $k\in\mathbb{N}$ and $(d_{1},\ldots,d_{k})\in\mathscr{D}^{k}$, if $(x_{1},\ldots,x_{n},c,d_{1},\ldots,d_{k})\in\mathsf{R}(n+k+1)$ then
@@
  1. i.
    The proof amounts to checking all the possible choices for $\mathfrak{F}^{\circ}_{n}(x_{1},\ldots,x_{n})$ and $c\in\mathscr{D}$. For example, if $\mathfrak{F}^{\circ}_{n}(x_{1},\ldots,x_{n})=\mathfrak{F}^{\circ}_{1}(2)$, then $c=im$ for some $m\in\mathbb{Z}$ with $|m|\geq 2$ or $c=m$ for some $m\in\mathbb{N}_{\geq 2}$ (see Figure 5); in either case, $\mathfrak{F}^{\circ}_{n+1}(x_{1},\ldots,x_{n},c)=\mathfrak{F}^{\circ}(c)$.
  2. ii.
    Assume $|c|\geq 3$. The previous part yields $\mathfrak{F}^{\circ}_{n+1}(x_{1},\ldots,x_{n},c)=\mathfrak{F}^{\circ}$ and, by (9),
@@
    (24) $=\mathfrak{F}^{\circ}_{k}(d_{1},\ldots,d_{k}).$
If $|c|=2$, then $\mathfrak{F}^{\circ}_{n+1}(x_{1},\ldots,x_{n},c)\neq\varnothing$ may take 4 possible forms. Checking case by case, we may see that $(x_{1},\ldots,x_{n},c,d_{1})\in\mathsf{R}(n+2)$ implies
(25) $\mathfrak{F}^{\circ}_{n+2}(x_{1},\ldots,x_{n},c,d_{1})=\mathfrak{F}^{\circ}_{1}(d_{1}).$
@@
$\iota[\mathfrak{F}^{\circ}_{n+1}(x_{1},\ldots,x_{n},2)]=\iota\left[\mathfrak{F}^{\circ}_{1}(2)\right]=\iota[\mathfrak{F}^{\circ}]\cap\left\{z\in\mathbb{C}:\operatorname{Re}(z)>-\frac{1}{2}\right\}.$
Thus, we get (25), because for all $d_{1}\in\mathscr{D}$ we have
$\tau_{d_{1}}[\mathfrak{F}^{\circ}]\cap\iota[\mathfrak{F}^{\circ}_{n+1}(x_{1},\ldots,x_{n},c)]=\begin{cases}\varnothing,&\text{ if }\operatorname{Re}(d_{1})<0,\\ \tau_{d_{1}}[\mathfrak{F}^{\circ}]\cap\iota[\mathfrak{F}^{\circ}]\neq\varnothing,&\text{ if }\operatorname{Re}(d_{1})\geq 0,\end{cases}$
(cfr. Figures 1, 5). We are done if $k=1$. If not, we proceed as in (24) and using (25). ∎
@@
If $a\in\{1+im,m+i\}$ for some $m\in\mathbb{Z}$, $|m|\geq 2$, then
@@
Proof.
The lemma follows from the definitions of $\mathsf{mir}_{1}$, $\mathsf{mir}_{2}$, and $S$. ∎
@@
Take $\mathbf{b}=(b_{n})_{n\geq 1}\in\Omega$. If $j\in\mathbb{N}$ satisfies $(b_{1},\ldots,b_{j})\in\mathsf{R}(j)$ and $(b_{1},\ldots,b_{j})\in\mathsf{Ir}(j+1)$, then
@@
This is a consequence of Lemma 3.5 and its proof. ∎
7.2. Proof of Proposition 4.1
We deduce Proposition 4.1 after showing that Procedure 4.3 gives the desired output. To this end, let $\mathbf{a}\in\mathsf{Ir}$ be arbitrary, and $(\mathbf{b}_{N})_{N\geq 1}$ and $\mathbf{b}$ as in Procedure 4.3. We will verify the next three assertions:
  1. A.
    The sequence $(\mathbf{b}_{N})_{N\geq 0}$ converges in the product topology to some $\mathbf{b}\in\mathscr{D}^{\mathbb{N}}$,
  2. B.
    $\mathbf{b}:=\displaystyle\lim_{N\to\infty}\mathbf{b}_{N}\in\overline{\mathsf{R}}$, and
  3. C.
    $\overline{\Lambda}(\mathbf{b})=\Lambda(\mathbf{a})$.
First, we show that the sequences $\mathbf{b}_{N}$ and $\mathbf{b}_{N+1}$ have increasingly longer regular prefixes in common; this will be proved below by induction. With that, Assertions A and B will follow. Our arguments will also imply that
$\Lambda(\mathbf{a})=[0;b_{N}(1),b_{N}(2),b_{N}(3),\ldots],$
which will lead us to conclude Assertion C.
Base of induction.
Let us start with $N=0$. Since every cylinder of level $1$ is regular, there is some $j_{0}\in\mathbb{N}$ such that
@@
Note that $\mathfrak{F}^{\circ}_{j_{0}}(a_{1},\ldots,a_{j_{0}})\neq\mathfrak{F}^{\circ}$, otherwise $(a_{1},\ldots,a_{j_{0}},a_{j_{0}+1})$ would be regular. By the proof of Lemma 3.5, we may conclude the existence of some $m\in\mathbb{Z}$ such that
@@
We verify that the sequence $\mathbf{b}_{1}=(b_{1}(j))_{j\geq 1}$ has the next properties:
(26) $(b_{1}(1),\ldots,b_{1}(j_{0}+1))\in\mathsf{R}(j_{0}+1),$
(27) $\Lambda(\mathbf{a})=[0;b_{1}(1),b_{1}(2),b_{1}(3),\ldots].$
Case $|m|\geq 3$. By Lemma 3.5, $\mathfrak{F}^{\circ}_{j_{0}+1}\left(a_{1},\ldots,a_{j_{0}},S(a_{j_{0}+1})\right)=\mathfrak{F}^{\circ}$ and (26) follows. Now we show (27). If $a_{j_{0}+1}=1+im$, then Lemma 7.3 and $[0;a_{j_{0}+2},a_{j_{0}+3},\ldots]_{\mathbb{C}}\in\mathfrak{F}$ give
@@
$=[1-\overline{a_{j_{0}+1}};-\overline{a_{j_{0}+2}},-\overline{a_{j_{0}+3}},\ldots]$
@@
$=[0;b_{1}(1),b_{1}(2),\ldots,b_{1}(j_{0}),b_{1}(j_{0}+1),b_{1}(j_{0}+2),b_{1}(j_{0}+3),\ldots].$
Similarly, if $a_{j_{0}+1}=m+i$, from Lemma 7.3 and $[0;a_{j_{0}+2},a_{j_{0}+3},\ldots]_{\mathbb{C}}\in\mathfrak{F}$ we get
@@
Since $z=\overline{z}+i$ when $\operatorname{Im}(z)=\frac{1}{2}$, we have
@@
$=[S\left(a_{j_{0}+1}\right);\mathsf{mir}_{1}(a_{j_{0}+1}),\mathsf{mir}_{1}(a_{j_{0}+2}),\mathsf{mir}_{1}(a_{j_{0}+3}),\ldots]$
and (26) follows.
Case $|m|=2$. Apply $\iota$ on the sets in Lemma 3.5 to obtain $a_{j_{0}+1}\in\{1+2i,1-2i,2+i,-2+i\}$. Suppose that $a_{j_{0}+1}=1+2i$, then either
@@
(the only sets in Lemma 3.5 leading to an irregular cylinder when the next partial quotient is $1+2i$ are $\mathfrak{F}^{\circ}_{1}(-1-i)$ and $\mathfrak{F}^{\circ}_{1}(-2)$). In either case, we have
$\mathfrak{F}^{\circ}_{j_{0}+1}\left(a_{1},\ldots,a_{j_{0}},S(a_{j_{0}+1})\right)=\mathfrak{F}^{\circ}_{j_{0}+1}(a_{1},\ldots,a_{j_{0}},2i)=\mathfrak{F}^{\circ}_{1}(2i)\neq\varnothing$
and we get (26). We show (27) as in the case $|m|\geq 3$. The possibilities $a_{j_{0}+1}\in\{1-2i,2+i,2-i\}$ are treated similarly.
Inductive step
Assume that for some $N\in\mathbb{N}$ we can perform $N$ iterations of the procedure and that for each $k\in\{1,\ldots,N\}$ we have
(30) $(b_{k}(1),\ldots,b_{k}(j_{k-1}+1))\in\mathsf{R}(j_{k-1}+1),$
(31) $\operatorname{Re}\left([a_{j_{k-1}+1};a_{j_{k-1}+2},\ldots]_{\mathbb{C}}\right)=\frac{1}{2}\;\text{ or }\;\operatorname{Im}\left([a_{j_{k-1}+1};a_{j_{k-1}+2},\ldots]_{\mathbb{C}}\right)=\frac{1}{2}\text{ and },$
(32) $\Lambda(\mathbf{a})=[0;b_{k}(1),b_{k}(2),b_{k}(3),\ldots].$
@@
Let us discuss the $(N+1)$th iteration. Suppose there is some $j_{N}\in\mathbb{N}$ satisfying
(33) $\left(b_{N}(1),\ldots,b_{N}(j_{N})\right)\in\mathsf{R}(j_{N})\;\text{ and }\;\left(b_{N}(1),\ldots,b_{N}(j_{N}+1)\right)\not\in\mathsf{R}(j_{N}+1).$
Assertions A, B, and C are trivially true if no such $j_{N}$ exists. We claim that
@@
(36) $\Lambda(\mathbf{a})=[0;b_{N+1}(1),b_{N+1}(2),b_{N+1}(3),\ldots].$
From (30), we deduce $j_{N}-j_{N-1}\geq 1$. We consider two cases: $j_{N}-j_{N-1}=1$ and $j_{N}-j_{N-1}\geq 2$.
Case $j_{N}=j_{N-1}+1$. Since $b_{N}(j_{N})=S(b_{N-1}(j_{N}))=S(b_{N-1}(j_{N-1}+1))$, there is some $m\in\mathbb{Z}$, $|m|\geq 2$, such that $b_{N}(j_{N})\in\{m,im,-m,-im\}$. If we had $|m|\geq 3$, then Lemma 7.1 would give
@@
which implies $(b_{N}(1),\ldots,b_{N}(j_{N}),b_{N}(j_{N}+1))\in\mathsf{R}(j_{N}+1)$, contradicting (33). Therefore, $|m|=2$ and $b_{N}(j_{N})\in\{-2,-2i,2,2i\}$. Let us further assume that $b_{N}(j_{N})=-2$, the other possibilities being treated similarly. Then, we have
@@
Suppose that $b_{N-1}(j_{N})=-2+i$, an analogous argument works for $-2-i$. Then, Condition (33) and Lemma 7.1 imply
(37) $\mathfrak{F}^{\circ}_{j_{N}}\left(b_{N}(1),\ldots,b_{N}(j_{N})\right)=\mathfrak{F}^{\circ}_{1}(-2)\quad\text{ and }\quad\operatorname{Re}\left(b_{N}(j_{N}+1)\right)\geq 1.$
To see the inequality in (37), note that $\mathfrak{F}^{\circ}_{1}(-2)=-\mathfrak{F}^{\circ}_{1}(2)$, $\iota[\mathfrak{F}^{\circ}_{1}(-2)]=-\iota[\mathfrak{F}(2)]$ and see Figure 5. We claim that
@@
Certainly, when $N=1$, we consider $\mathsf{mir}$ to be the identity map on $\mathbb{C}$. For $N>1$, the induction hypothesis gives $j_{N-1}\geq j_{N-2}+1$, so $j_{N-1}+n\geq j_{N-2}+n+1$ for all $n\in\mathbb{N}$ and (38) follows from Remark 7.4.
Next, we separately consider every possibility for $\mathsf{mir}$ in (38). Recall that
@@
  1. i.
    When $\mathsf{mir}$ is the identity map, we have $a_{j_{N}}=-2+i$ and $\operatorname{Re}(a_{j_{N}+1})\geq 1$. Hence, the set $\mathfrak{F}_{j_{N}}(a_{1},\ldots,a_{j_{N}})\in\mathsf{Ir}(j_{N})$ is of one of the forms depicted in Figure 8. To check this observation, we determine inductively on $n$ the possible forms of $\mathfrak{F}_{n}(\mathbf{c})$, $\mathbf{c}\in\mathsf{Ir}(n)$. We consider each form independently.
    1. (i.a).
      Assume that $\mathfrak{F}_{j_{N}}(a_{1},\ldots,a_{j_{N}})=\{[0;a_{j_{N}+1},a_{j_{N}+2},\ldots]_{\mathbb{C}}\}=\{\zeta_{4}\}$. This happens, for example, when $(a_{1},\ldots,a_{j_{N}})=(2i,-2+i,1+2i,-2+i)$. We conclude (35) from $\operatorname{Re}(\zeta_{4}^{-1})=\frac{1}{2}$. Moreover, we have
@@
      $[b_{N}(j_{N});b_{N}(j_{N}+1),\ldots]=[b_{N+1}(j_{N});b_{N+1}(j_{N}+1),\ldots].$
      Since $b_{N}(k)=b_{N+1}(k)$ for all $k\in\{1,\ldots,j_{N}-1\}$, Equation (36) follows.
    2. (i.b).
      Suppose that $\mathfrak{F}_{j_{N}}(a_{1},\ldots,a_{j_{N}})$ is the arch of $C(1-i)$ going from $-\zeta_{1}=\frac{1}{2}-i\alpha$ (excluded) to $\zeta_{4}$ (included). An example of this case is $(a_{1},\ldots,a_{j_{N}})=(-2,1+3i,-2+i)$. Then,
@@
      so $\zeta_{2}=[0;a_{j_{N+1}},a_{j_{N+2}},\ldots]$ and we argue as in Case i.(i.a).
    3. (i.c).
      When $\mathfrak{F}_{j_{N}}(a_{1},\ldots,a_{j_{N}})$ is the arch of $C(1)$ from $\overline{\zeta_{4}}$ (excluded) to $\zeta_{4}$ (included), we argue as in Case i.(i.b). An instance of this case is $(a_{1},\ldots,a_{j_{N}})=(2i,-3+i,1+2i,-2+i)$.
    4. (i.d).
@@
      and $\zeta_{2}=[0;a_{j_{N}+1},a_{j_{N}+2},\ldots]_{\mathbb{C}}$. Now, we proceed as in Case i.(i.a). An instance of this case is $(a_{1},\ldots,a_{j_{N}})=(2i,-2+i)$.
@@
  2. ii.
    If $\mathsf{mir}=\mathsf{mir}_{1}$, then $a_{j_{N}}=-2-i$ but (31) fails for $k=N$.
  3. iii.
    When $\mathsf{mir}=\mathsf{mir}_{2}$, we argue as when $\mathsf{mir}$ is the identity map so we omit some details. We have $a_{j_{N}}=2+i$, $\operatorname{Re}(a_{j_{N}+1})\leq-1$ and, by (31) for $k=N$,
    (39) $\operatorname{Im}([a_{j_{N}};a_{j_{N}+1},a_{j_{N}+2},\ldots]_{\mathbb{C}})=\frac{1}{2}.$
@@
    1. (iii.a).
      If $\mathfrak{F}_{j_{N}}(a_{1},\ldots,a_{j_{N}})=\{\zeta_{2}\}=\{[0;a_{j_{N}+1},a_{j_{N}+2},\ldots]_{\mathbb{C}}\}$, then $[a_{j_{N}};a_{j_{N}+1},\ldots]_{\mathbb{C}}=-2+i+\zeta_{2}$, contradicting (39).
    2. (iii.b).
      If $\mathfrak{F}_{j_{N}}(a_{1},\ldots,a_{j_{N}})$ were the arch of $C(-1-i)$ going from $\zeta_{3}$ (included) to $\zeta_{2}$ (included), by (31) we would obtain $\zeta_{3}=[0;a_{j_{N}+1},a_{j_{N}+2},\ldots]_{\mathbb{C}}$ so $a_{j_{N}+1}=2i$, but this contradicts $\operatorname{Re}(a_{j_{N}+1})\leq-1$.
    3. (iii.c).
      The impossibility of $\mathfrak{F}_{j_{N}}(a_{1},\ldots,a_{j_{N}})$ being the arch of $C(-1)$ from $\zeta_{2}$ (included) to $\overline{\zeta_{2}}$ (excluded) is shown as in Case iii.(iii.b).
    4. (iii.d).